# Knowledge Transfer from Simple to Complex: A Safe and Efficient Reinforcement Learning Framework for Autonomous Driving Decision-Making

### v3

@@ Abstract
A safe and efficient decision-making system is crucial for autonomous vehicles. However, the complexity of driving environments limits the effectiveness of many rule-based and machine learning approaches. Reinforcement Learning, with its robust self-learning capabilities and environmental adaptability, offers a promising solution to these challenges. Nevertheless, safety and efficiency concerns during training hinder its widespread application. To address these concerns, we propose a novel RL framework, Simple to Complex Collaborative Decision (S2CD). First, we rapidly train the teacher model in a lightweight simulation environment. In the more complex and realistic environment, the teacher intervenes when the student agent exhibits suboptimal behavior by assessing actions’ value to avert dangers. We also introduce an RL algorithm called Adaptive Clipping Proximal Policy Optimization (ACPPO), which combines samples from both teacher and student policies and employs dynamic clipping strategies based on sample importance. This approach improves sample efficiency while effectively alleviating data imbalance. Additionally, we employ the Kullback-Leibler divergence as a policy constraint, transforming it into an unconstrained problem with the Lagrangian method to accelerate the student’s learning. Finally, a gradual weaning strategy ensures that the student learns to explore independently over time, overcoming the teacher’s limitations and maximizing performance. Simulation experiments in highway lane-change scenarios show that the S2CD framework enhances learning efficiency, reduces training costs, and significantly improves safety compared to state-of-the-art algorithms. This framework also ensures effective knowledge transfer between teacher and student models, even with a suboptimal teacher, the student achieves superior performance, demonstrating the robustness and effectiveness of S2CD.
@@ 1 Introduction
Due to the complexity and unpredictability of driving environments, approximately 94% of traffic accidents are related to suboptimal decisions made by human drivers (Alvaro et al., 2018). Suboptimal driving decisions not only compromise traffic safety but also reduce overall traffic efficiency (Lee et al., 2004). Hopefully, the rapid advancement of autonomous driving technology has attracted significant attention due to its potential to improve both traffic safety and efficiency. Autonomous driving systems generally consist of a perception layer, a decision-making layer, and a planning and control layer, in which the decision-making layer commonly referred to as the "brain" of the system (Chen et al., 2017). Ensuring the safety and accuracy of decision-making algorithms is therefore essential for the overall performance of autonomous driving systems (Li et al., 2017). Currently, decision-making methods for autonomous vehicles are generally classified into two categories: knowledge-driven and data-driven. Knowledge-driven methods, such as Hierarchical State Machines (HSM) (Patz et al., 2008), Expert Systems (ES) (Ferguson et al., 2008), and Finite State Machines (FSM) (Buehler et al., 2009), are known for their rigorous logic and high interpretability (Bae et al., 2020). However, these methods are limited by their strong reliance on existing knowledge, which restricts their ability to handle novel or unexpected scenarios (Bianco-Vega et al., 2005; Mozina et al., 2008). In contrast, data-driven methods such as End-to-End (E2E) Learning (Bojarski et al., 2016), Imitation Learning (IL) (Hawke et al., 2020), and Reinforcement Learning (RL) (Sallab et al., 2017; Zhang et al., 2022c) have gained prominence due to their robust self-learning capabilities and adaptability to dynamic environments. RL algorithms, in particular, optimize driving strategies through real-time interactions with the environment, making them highly effective in complex scenarios (Zhu and Zhao, 2021). While RL has demonstrated success in fields such as game artificial intelligence (AI) (Zhang et al., 2022a; Ye et al., 2020), financial transactions (Tsantekidis et al., 2023; Pendharkar and Cusatis, 2018), and energy management (Yu et al., 2019; Kuznetsova et al., 2013), its application in safety-critical areas like autonomous driving and robotic control faces significant challenges due to safety risks and training inefficiency (Liu et al., 2022; Ou et al., 2024).
@@
In summary, traditional RL algorithms involve high training costs but fail to guarantee the safety of agents during training. While standard TSF methods help student agents learn safely and efficiently, they come with significant costs in developing a high-performance teacher, and the student’s ultimate performance is constrained by the teacher’s abilities. Furthermore, these methods cannot optimize the policy by simultaneously utilizing samples generated by both the teacher and student models. To address these limitations, this paper introduces a novel TSF framework called Simple to Complex Collaborative Decision (S2CD), which enhances knowledge transfer between teacher and student. Unlike other TSFs, the S2CD framework accelerates training by first training the teacher model in a lightweight simulation environment with similar scenarios, then guiding the student in more complex simulations or real-world settings. We also theoretically demonstrated that teacher guidance improves both the performance and safety ceilings of the student during training. Additionally, we enhance the Proximal Policy Optimization (PPO) algorithm and propose Adaptive Clipping PPO (ACPPO), which utilizes data from both teacher and student models, significantly improving sample efficiency and alleviating data imbalance. The algorithm adaptively adjusts the clipping factor based on sample importance, boosting learning efficiency. During the policy optimization process, the neural network is updated by constraining the Kullback-Leibler (KL) divergence between the teacher’s and student’s policies, facilitating the rapid convergence of the student policy to the teacher’s. Finally, S2CD employs a weaning strategy that allows the student agent to independently explore the environment in later training stages, thus overcoming the teacher’s limitations. The main contributions of this paper are summarized as follows:
@@
  • This paper proposes the ACPPO algorithm, integrated within the S2CD framework, which simultaneously utilizes data from both teacher and student models while adaptively adjusting the clipping factor based on sample importance, thus enhancing sample efficiency and alleviating data imbalance.
  • This paper introduces the use of the KL divergence between the teacher’s and student’s policies as a constraint during policy updates, transforming it into an unconstrained problem using the Lagrangian method to facilitate the rapid convergence of the student’s policy to the teacher’s.
  • This paper investigates intervention and weaning strategies within the S2CD framework, where teacher interventions are triggered based on action values, and the frequency of interventions is gradually reduced to promote the student’s independent exploration, mitigating the limitations imposed by suboptimal teacher performance.
Section 2 reviews related work and analyzes the strengths and weaknesses of these studies to underscore the value of our research. Section 3 elaborates on the theoretical foundation of the PPO algorithm and the principles of the ACPPO algorithm before introducing the S2CD decision-making framework. This framework comprises an offline learning-based teacher training module, an ACPPO-based high-level decision-making layer, a cubic spline interpolation-based path planning layer, and a PID algorithm-based control layer. Section 4 provides a detailed explanation of the differences between the Highway-Env and Carla simulation environments, as well as the implementation details of the simulation experiments, including scenario specification, scene modeling, and simulation settings. Section 5 presents the simulation results of the S2CD framework and compares them with other mainstream algorithms. Section 6 concludes the paper and then explores the study’s limitations and directions for future research.
@@ 2.1 Knowledge Transfer
For example, Qiao et al. (2018) and Song et al. (2021) employed Curriculum Learning (CL) strategies to achieve notable performance in urban intersection and overtaking tasks, respectively. CL enhances model learning effectiveness by gradually increasing task difficulty, instead of requiring models to tackle complex tasks directly (Bengio et al., 2009). However, CL requires meticulously designed multi-level tasks. Furthermore, the progression from simpler to more complex tasks can increase training time and computational resource demands. Transfer Learning (TL) is another knowledge transfer technique that effectively reduces the training costs of the target task by applying knowledge acquired from the source task to the target task. For instance, Shu et al. (2021) combined RL with TL to transfer decision-making knowledge from one driving task to another in an intersection environment. Similarly, Akhauri et al. (2020) utilized simulated accident data for TL, enabling the model to generalize more rapidly and effectively to real-world scenarios. However, if the source and target tasks are not sufficiently related, the performance of TL may deteriorate, potentially resulting in "negative transfer" (Zhang et al., 2022b). Another approach to knowledge transfer is the use of Digital Twins (DT). By creating an accurate digital replica of a physical entity or system, we can map data from the physical world to the virtual world (Niaz et al., 2021), enabling real-time monitoring, simulation, analysis, and optimization of physical entities. Voogd et al. (2023) employed DT technology to combine virtual and real-world data on vehicle dynamics and traffic scenarios, effectively bridging the gap between simulation and reality. Wang et al. (2024) designed and implemented an E2E DT system for autonomous driving. This system captures real-world traffic information, processes it in the cloud to create a digital twin model, and then utilizes this model to provide route planning services for autonomous vehicles. Despite DT technology effectively bridging the gap between simulation training and real-world applications, accurate simulation processes remain indispensable. Designing simulation programs that accurately mirror or fully replicate real-world conditions is both challenging and costly (Yun and Park, 2021). Moreover, pre-training a high-performance model in a high-fidelity simulation environment is resource-intensive.
@@ 2.2 Safe RL
For instance, methods like PPO-Lag (Stooke et al., 2020) and SAC-Lag (Ha et al., 2021) utilize the Lagrangian method, which simultaneously updates both the policy and the Lagrange multiplier through primal-dual optimization, thereby transforming the safety-constrained optimization problem into an unconstrained one. Peng et al. (2022a) proposed the Separated Proportional Integral Lagrangian (SPIL) algorithm, which mitigates oscillations and reduces conservatism in RL policies during car-following scenarios. Zhang et al. (2024) introduced cognitive uncertainty as a constraint in Lagrangian Safe RL algorithms, aiming to enhance both exploration and safety performance in autonomous vehicles. While the Lagrangian method is effective in handling safety constraints, it lacks a clear mechanism to prevent hazardous events (Li et al., 2022b). Thus, agents may still engage in unsafe exploration during training. Consequently, many researchers have applied rule-based constraints (Cao et al., 2022; Likmeta et al., 2020) or risk estimation methods (Mo et al., 2021; Li et al., 2022a) to Safe RL in autonomous driving, aiming to ensure a minimum level of policy safety performance. Although these methods can prevent a significant number of unsafe actions, their effectiveness heavily depends on the quality of the rule design or risk assessment models. To further reduce unsafe events, researchers have shifted their focus toward Safe RL approaches based on the TSF. Within the TSF, teacher policies can be classified into human expert strategies and well-trained neural network strategies. For example, Wu et al. (2023) and Huang et al. (2024) employed human experts to mentor AI agents. During the student agent’s interaction with the environment, the human expert intervenes in dangerous situations and demonstrates the correct actions, thereby ensuring driving safety throughout the student’s training process. While Safe RL methods involving human expert intervention significantly improve the safety of the exploration process, they incur high labor costs. Moreover, the teaching performance of human experts is highly influenced by subjective factors, leading to instability in the final model performance. To reduce labor costs, many researchers have employed well-trained neural network models as "teachers" to guide student agents during the learning process. For instance, Peng et al. (2022b) and Xue et al. (2023) employed RL algorithms to obtain a well-trained teacher model, when the conditions defined by a switching function, which determines whether to intervene, are met, the teacher takes over the student’s actions to demonstrate correct operations and avoid dangerous situations. Zhou et al. (2024) employed the SOAR cognitive architecture as a teacher to guide the student agent toward efficient and safe learning, while also offering a certain degree of interpretability.
@@ 3 Theoretical Background
In this section, we provide a comprehensive explanation of the technical background underlying the S2CD framework. First, we present the theoretical foundations of the PPO algorithm. Subsequently, we refine the clipping strategy of the PPO algorithm, introducing the ACPPO algorithm. Lastly, we incorporate ACPPO into the S2CD framework and offer a detailed description of the components within the S2CD framework.
@@ 3.1 Proximal Policy Optimization Algorithm
DRL achieves its goals through interactions between an agent and its environment, and it can be broadly categorized into two main approaches: value-based methods and policy-based methods. The fundamental principle of value-based methods is to learn a value function that estimates the long-term returns for each state or state-action pair. The optimal action is then selected based on this value function. In contrast, policy-based methods directly optimize a policy that maps states to actions, making decisions based on the probabilistic distribution of this policy. Although value-based methods perform adequately in simpler tasks, their efficiency decreases significantly as the complexity of the state or action space increases. Conversely, policy-based DRL methods offer a well-defined optimization path that ensures convergence to a stable policy while enhancing exploration through the learning of stochastic policies. In high-dimensional and complex task environments, policy-based methods are typically more appropriate and offer superior scalability. Therefore, this paper will focus on policy-based DRL methods.
The objective of DRL is to identify a policy $\pi(a|s)$ that maximizes the cumulative discounted reward across all time steps, defined as:
@@
where $\mathbb{E}$ is the expectation operator for a finite batch of samples, $\pi_{\theta}$ denotes the current policy, with $\theta$ representing its parameters. The state value function $V^{\pi}(s_{t})=\mathbb{E}_{\pi}\left[\sum_{k=0}^{\infty}\gamma^{k}r(s_{t+k},a_{t+k})\middle|s_{t}\right]$ represents the expected return when starting from state $s_{t}$ under policy $\pi_{\theta}$.
@@
$\begin{split}J(\theta)&=\mathbb{E}_{s_{0}}\left[V^{\pi_{\theta}}(s_{0})\right]\\ &=\mathbb{E}_{\pi_{\theta^{\prime}}}\left[\sum_{t=0}^{\infty}\gamma^{t}V^{\pi_{\theta}}(s_{t})-\sum_{t=1}^{\infty}\gamma^{t}V^{\pi_{\theta}}(s_{t})\right]\\ &=-\mathbb{E}_{\pi_{\theta^{\prime}}}\left[\sum_{t=0}^{\infty}\gamma^{t}\left(\gamma V^{\pi_{\theta}}(s_{t+1})-V^{\pi_{\theta}}(s_{t})\right)\right]\end{split}$ (3)
According to above equations and Paper Kakade and Langford (2002), we can deduce the gap between the objective function of the new policy $\pi_{\theta^{\prime}}$ and the current policy $\pi_{\theta}$:
@@
$\centering A^{\pi_{\theta}}(s_{t},a_{t})&=Q^{\pi_{\theta}}(s_{t},a_{t})-V^{\pi_{\theta}}(s_{t})\\ &=r(s_{t},a_{t})+\gamma V^{\pi_{\theta}}(s_{t+1})-V^{\pi_{\theta}}(s_{t})\end{split}$ (5)
@@
$\begin{split}J(\theta^{\prime})-J(\theta)&=\mathbb{E}_{\pi_{\theta^{\prime}}}\left[\sum_{t=0}^{\infty}\gamma^{t}A^{\pi_{\theta}}(s_{t},a_{t})\right]\\ &=\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}_{s_{t}\sim P_{t}^{\pi_{\theta^{\prime}}}}\mathbb{E}_{a_{t}\sim\pi_{\theta^{\prime}}(\cdot|s_{t})}\left[A^{\pi_{\theta}}(s_{t},a_{t})\right]\end{split}$ (6)
@@
$\begin{split}J(\theta^{\prime})-J(\theta)&=\dfrac{1}{1-\gamma}\sum_{s}\nu^{\pi_{\theta^{\prime}}}(s)\sum_{a}\pi_{\theta^{\prime}}(a|s)A^{\pi_{\theta}}(s,a)\\ &=\dfrac{1}{1-\gamma}\mathbb{E}_{s\sim\nu^{\pi_{\theta^{\prime}}}}\mathbb{E}_{a\sim\pi_{\theta^{\prime}}(\cdot|s_{t})}[A^{\pi_{\theta}}(s,a)]\end{split}$ (7)

### v4

@@ Abstract
A safe and efficient decision-making system is crucial for autonomous vehicles. However, the complexity of driving environments limits the effectiveness of many rule-based and machine learning approaches. Reinforcement Learning (RL), with its robust self-learning capabilities and environmental adaptability, offers a promising solution to these challenges. Nevertheless, safety and efficiency concerns during training hinder its widespread application. To address these concerns, we propose a novel RL framework, Simple to Complex Collaborative Decision (S2CD). First, we rapidly train the teacher model in a lightweight simulation environment. In the more complex and realistic environment, teacher intervenes when the student agent exhibits suboptimal behavior by assessing actions’ value to avert dangers. We also introduce an RL algorithm called Adaptive Clipping Proximal Policy Optimization Plus, which combines samples from both teacher and student policies and employs dynamic clipping strategies based on sample importance. This approach improves sample efficiency while effectively alleviating data imbalance. Additionally, we employ the Kullback-Leibler divergence as a policy constraint, transforming it into an unconstrained problem with the Lagrangian method to accelerate the student’s learning. Finally, a gradual weaning strategy ensures that the student learns to explore independently over time, overcoming the teacher’s limitations and maximizing performance. Simulation experiments in highway lane-change scenarios show that the S2CD framework enhances learning efficiency, reduces training costs, and significantly improves safety compared to state-of-the-art algorithms. This framework also ensures effective knowledge transfer between teacher and student models, even with suboptimal teachers, the student achieves superior performance, demonstrating the robustness and effectiveness of S2CD.
@@ 1 Introduction
Given the unpredictable and complex nature of driving environments, nearly 94% of traffic incidents are related to poor decision-making by human drivers (Alvaro et al., 2018). Suboptimal driving decisions negatively impact both the safety and efficiency of traffic flow (Lee et al., 2004). With the rapid progress in autonomous driving technology, considerable interest has been drawn to its promise for enhancing both traffic safety and efficiency. Autonomous vehicle systems generally consist of a perception layer, a decision-making layer, and a planning and control layer, in which the decision-making layer commonly referred to as the "brain" of the system (Chen et al., 2017). Therefore, the performance of decision-making algorithms directly relates to the effectiveness of intelligent vehicle technologies (Li et al., 2017). Currently, decision-making methods for autonomous vehicles are generally classified into two categories: knowledge-driven and data-driven. Knowledge-driven methods, such as Hierarchical State Machines (HSM) (Patz et al., 2008), Expert Systems (ES) (Ferguson et al., 2008), and Finite State Machines (FSM) (Buehler et al., 2009), are known for their rigorous logic and high interpretability (Bae et al., 2020). However, these methods are limited by their strong reliance on existing knowledge, which restricts their ability to handle novel or unexpected scenarios (Bianco-Vega et al., 2005; Mozina et al., 2008). In contrast, data-driven methods such as End-to-End (E2E) Learning (Bojarski et al., 2016), Imitation Learning (IL) (Hawke et al., 2020), and Reinforcement Learning (RL) (Sallab et al., 2017; Zhang et al., 2022c) have gained prominence due to their robust self-learning capabilities and adaptability to dynamic environments. RL algorithms, in particular, optimize driving strategies through real-time interactions with the environment, making them highly effective in complex scenarios (Zhu and Zhao, 2021). While RL has demonstrated success in fields such as game artificial intelligence (AI) (Zhang et al., 2022a; Ye et al., 2020), financial transactions (Tsantekidis et al., 2023; Pendharkar and Cusatis, 2018), and energy management (Yu et al., 2019; Kuznetsova et al., 2013), its application in safety-critical areas like autonomous driving and robotic control faces significant challenges due to safety risks and training inefficiency (Liu et al., 2022; Ou et al., 2024).
@@
In summary, traditional RL algorithms involve high training costs and fail to ensure the safety of agents. While standard TSF methods help student agents learn safely and efficiently, they come with significant costs in developing a high-performance teacher, and the student’s ultimate performance is constrained by the teacher’s abilities. Furthermore, these methods cannot optimize the policy by simultaneously utilizing samples generated by both the teacher and student models. To tackle these shortcomings, this paper introduces an innovative TSF framework called Simple to Complex Collaborative Decision (S2CD), which enhances knowledge transfer between teacher and student. Unlike other TSFs, the S2CD framework accelerates training by first training the teacher in a simple simulation environment with similar scenarios, then guiding the student in more complex simulations or real-world settings. We also theoretically demonstrated that teacher guidance improves both the performance and safety ceilings of the student during training. Additionally, we propose the Adaptive Clipping Proximal Policy Optimization Plus (ACPPO+), which builds upon our previous work, ACPPO (Zhou et al., 2024), by leveraging data from both teacher and student models to further improve sample efficiency and mitigate data imbalance. The algorithm adaptively adjusts the clipping factor based on sample importance, boosting learning efficiency. During the policy optimization process, the neural network is updated by constraining the Kullback-Leibler (KL) divergence between the teacher’s and student’s policies, facilitating the rapid convergence of the student policy to the teacher’s. Finally, S2CD employs a weaning strategy that allows the student agent to independently explore the environment in later training stages, thus overcoming the teacher’s limitations. The main contributions of this paper are summarized as follows:
@@
  • This paper proposes the ACPPO+ algorithm, integrated within the S2CD framework, which simultaneously utilizes data from both teacher and student models while adaptively adjusting the clipping factor according to sample importance, thus enhancing sample efficiency and mitigating data imbalance.
  • This paper introduces the use of the KL divergence between the teacher’s and student’s policies as a constraint during policy updates, and transforms it into an unconstrained problem using the Lagrangian method to facilitate the rapid convergence of the student’s policy to the teacher’s.
  • This paper relies on action values to trigger teacher policy intervention and gradually reduces the probability of intervention through a weaning mechanism, encouraging independent exploration by the student and mitigating the limitations imposed by suboptimal teacher performance.
Section 2 reviews related work and analyzes the strengths and weaknesses of these studies to underscore the value of our research. Section 3 provides a detailed explanation of the theoretical background of the algorithms, followed by an introduction to the S2CD decision-making framework. Section 4 describes the differences between the Highway-Env and Carla simulation environments and provides details on the implementation of the experiments. Section 5 shows the results of the simulation experiments and compares them with state-of-the-art (SOTA) algorithms. Section 6 concludes the paper, explores the limitations of the study, and suggests directions for future research.
@@ 2.1 Knowledge Transfer
For example, Qiao et al. (2018) and Song et al. (2021) employed Curriculum Learning (CL) strategies to achieve notable performance in urban intersection and overtaking tasks, respectively. CL enhances model learning effectiveness by gradually increasing task difficulty, instead of requiring models to tackle complex tasks directly (Bengio et al., 2009). However, CL requires meticulously designed multi-level tasks. Furthermore, the progression from simpler to more complex tasks can increase training time and computational resource demands. Transfer Learning (TL) is another knowledge transfer technique that effectively reduces the training costs of the target task by applying knowledge acquired from the source task to the target task. For instance, Shu et al. (2021) combined RL with TL to transfer decision-making knowledge from one driving task to another in an intersection environment. Similarly, Akhauri et al. (2020) utilized simulated accident data for TL, enabling the model to generalize more rapidly and effectively to real-world scenarios. However, if the target and source tasks are not adequately related, the performance of TL may deteriorate, potentially resulting in "negative transfer" (Zhang et al., 2022b). Another approach to knowledge transfer is the use of Digital Twins (DT). By creating an accurate digital replica of a physical entity or system, we can map data from the physical world to the virtual world (Niaz et al., 2021), enabling real-time monitoring, simulation, analysis, and optimization of physical entities. Voogd et al. (2023) employed DT technology to combine virtual and real-world data on vehicle dynamics and traffic scenarios, effectively bridging the gap between simulation and reality. Wang et al. (2024) designed and implemented an E2E DT system for autonomous driving. This system captures real-world traffic information, processes it in the cloud to create a digital twin model, and then utilizes this model to provide route planning services for autonomous vehicles. Despite DT technology effectively bridging the gap between simulation training and real-world applications, accurate simulation processes remain indispensable. Designing simulation programs that accurately mirror or fully replicate real-world conditions is both challenging and costly (Yun and Park, 2021). Moreover, pre-training a high-performance model in a high-fidelity simulation environment is resource-intensive.
@@ 2.2 Safe RL
For instance, methods like SAC-Lag (Ha et al., 2021) and PPO-Lag (Stooke et al., 2020) utilize the Lagrangian method, which simultaneously updates both the policy and the Lagrange multiplier through primal-dual optimization, thereby transforming the safety-constrained optimization problem into an unconstrained one. Peng et al. (2022a) proposed the Separated Proportional Integral Lagrangian (SPIL) algorithm, which mitigates oscillations and reduces conservatism in RL policies during car-following scenarios. Zhang et al. (2024) introduced cognitive uncertainty as a constraint in Lagrangian Safe RL algorithms, aiming to enhance both exploration and safety performance in autonomous vehicles. While the Lagrangian method is effective in handling safety constraints, it lacks a clear mechanism to prevent hazardous events (Li et al., 2022b). Thus, agents may still engage in unsafe exploration during training. Consequently, many researchers have applied rule-based constraints (Cao et al., 2022; Likmeta et al., 2020) or risk estimation methods (Mo et al., 2021; Li et al., 2022a) to Safe RL in autonomous driving, aiming to ensure a minimum level of policy safety performance. Although these methods can prevent a significant number of unsafe actions, their effectiveness heavily depends on the quality of the rule design or risk assessment models. To further reduce unsafe events, researchers have shifted their focus toward Safe RL approaches based on the TSF. Within the TSF, teacher policies can be classified into human expert strategies and well-trained neural network strategies. For example, Wu et al. (2023) and Huang et al. (2024) employed human experts to mentor AI agents. During the student agent’s interaction with the environment, the human expert intervenes in dangerous situations and demonstrates the correct actions, thereby ensuring driving safety throughout the student’s training process. While Safe RL methods involving human expert intervention significantly improve the safety of the exploration process, they incur high labor costs. Moreover, the teaching performance of human experts is highly influenced by subjective factors, leading to instability in the final model performance. To reduce labor costs, many researchers have employed well-trained neural network models as "teachers" to guide student agents during the learning process. For instance, Peng et al. (2022b) and Xue et al. (2023) employed RL algorithms to obtain a well-trained teacher model, when the conditions defined by a switching function, which determines whether to intervene, are met, the teacher takes over the student’s actions to demonstrate correct operations and avoid dangerous situations. Zhou et al. (2024) employed the SOAR cognitive architecture as a teacher to guide the student agent toward efficient and safe learning, while also offering a certain degree of interpretability.
@@ 3 Theoretical Background
In this section, we offer a comprehensive explanation of the technical background underlying the S2CD framework. First, we present the principle of the PPO algorithm. Subsequently, we introduce the ACPPO+ algorithm. Lastly, we offer a detailed description of the components within the S2CD framework.
@@ 3.1 Proximal Policy Optimization Algorithm
DRL methods are generally categorized into two types: value-based and policy-based methods. The core principle of value-based methods is to learn a value function that estimates the long-term return for each state or state-action pair, which is then used to guide action selection. In contrast, policy-based methods directly optimize a policy that maps states to actions, enabling decision-making based on this learned policy. Although value-based RL methods perform well in many simple tasks, their efficiency and accuracy significantly decrease as the state or action space becomes more complex. Conversely, policy-based methods provide a clear optimization pathway, ensuring convergence and enhancing exploration through stochastic policies. For complex tasks, policy-based methods are generally more suitable and offer greater scalability. Therefore, this study adopts policy-based RL methods.
The objective of RL is to identify a policy $\pi(a|s)$ that maximizes the cumulative discounted reward across all time steps, defined as:
@@
where $\mathbb{E}$ is the expectation for a batch of samples, $\pi_{\theta}$ denotes the current policy, with $\theta$ representing its parameters. The state value function $V^{\pi}(s_{t})=\mathbb{E}_{\pi}\left[\sum_{k=0}^{\infty}\gamma^{k}r(s_{t+k},a_{t+k})\middle|s_{t}\right]$ represents the expected return when starting from state $s_{t}$ under policy $\pi_{\theta}$.
@@
$\begin{split}J(\theta)=\mathbb{E}_{s_{0}}\left[V^{\pi_{\theta}}(s_{0})\right]=\mathbb{E}_{\pi_{\theta^{\prime}}}\left[\sum_{t=0}^{\infty}\gamma^{t}V^{\pi_{\theta}}(s_{t})-\sum_{t=1}^{\infty}\gamma^{t}V^{\pi_{\theta}}(s_{t})\right]=-\mathbb{E}_{\pi_{\theta^{\prime}}}\left[\sum_{t=0}^{\infty}\gamma^{t}\left(\gamma V^{\pi_{\theta}}(s_{t+1})-V^{\pi_{\theta}}(s_{t})\right)\right]\end{split}$ (3)
Based on the Eq. 2Eq. 3 and Paper Kakade and Langford (2002), we have:
@@
$\centering A^{\pi_{\theta}}(s_{t},a_{t})=Q^{\pi_{\theta}}(s_{t},a_{t})-V^{\pi_{\theta}}(s_{t})=r(s_{t},a_{t})+\gamma V^{\pi_{\theta}}(s_{t+1})-V^{\pi_{\theta}}(s_{t})\end{split}$ (5)
@@
$\begin{split}J(\theta^{\prime})-J(\theta)=\mathbb{E}_{\pi_{\theta^{\prime}}}\left[\sum_{t=0}^{\infty}\gamma^{t}A^{\pi_{\theta}}(s_{t},a_{t})\right]=\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}_{s_{t}\sim P_{t}^{\pi_{\theta^{\prime}}}}\mathbb{E}_{a_{t}\sim\pi_{\theta^{\prime}}(\cdot|s_{t})}\left[A^{\pi_{\theta}}(s_{t},a_{t})\right]\end{split}$ (6)
@@
$\begin{split}J(\theta^{\prime})-J(\theta)=\dfrac{1}{1-\gamma}\sum_{s}\nu^{\pi_{\theta^{\prime}}}(s)\sum_{a}\pi_{\theta^{\prime}}(a|s)A^{\pi_{\theta}}(s,a)=\dfrac{1}{1-\gamma}\mathbb{E}_{s\sim\nu^{\pi_{\theta^{\prime}}}}\mathbb{E}_{a\sim\pi_{\theta^{\prime}}(\cdot|s_{t})}[A^{\pi_{\theta}}(s,a)]\end{split}$ (7)